OpenStudy (anonymous):

Hi, Q: A solution of volume 0.500 contains 1.68g NH3 and 4.05g (NH4)2SO4. find the pH of the solution. I got 9.76 but that's incorrect and I have no idea what to do here's my work so I converted NH3 and (NH4)2SO4 into moles and divided them by 0.500L so the concentration of [NH3]=0.19718M and [(NH4)2SO4]=0.0613M then I used an ice table NH3 + H2O <--> NH4 + OH kb=1.8x10^-5 0.19718 0.061289 -x +x +x 0.19718-x 0.061289+x x kb=x(0.061289+x)/0.19718-x kb<<<<0.061289 and 0.19718

OpenStudy (anonymous):

kb=0.061289x/0.19718 x=1.0x10^-9 x is the concentration of OH pH=-log(1.0x10^-14/1.0x10^-9)=9.76